Draw the interference pattern for a double-slit interference situation on a screen. Label the constructive and destructive interference regions clearly with words and show the representation clearly. Indicate where the central maximum is. If a pane of glass is inserted in the gap between the slits and screen, what is the effect on the interference pattern? Draw the new pattern next to the original and explain any changes in the positionings of the pattern. Use words to explain as well as equations.
Answer #1

destructive interference constructive interference y d D Central Maxima 8

Glass D Ax sho D Central Maxima 8

There will be shift in the fringes as well as the position of the central maxima will get shifted by an amount of ∆y that is represented in the diagram. (Sry i hv written tht ∆x but its actually ∆y)

To calculate the change in the fringe pattern distance we have to calculate ∆x for the new arrangement:-

∆x= dsinø + t(m-1). Where t= thickness, m= refractive index.

=> ∆x= dy/D + t(m-1)

Clearly ∆y= Dt(m-1)/d. (Ans)

Except shifting of fringes everything remains unchanged.